# tenSVD algorithm for compression

### v1

@@ 3.2 Memory Cost
When lossy compression is the strategy, to reduce memory cost, the parameters $(r_{1},r_{2},\dots,r_{N})$ are fixed lower than $(R_{1},R_{2},\dots,R_{N})$. In the case where the elements are reordered into a higher-order tensor, it is not clear whether the memory cost will be higher. The storage requirements are given by the sum of the memory costs for each component matrix, represented as $\sum_{n=1}^{N}I_{n}\times r_{n}$, and the core tensor, represented as $\prod_{n=1}^{N}r_{n}$. If the elements are rearranged into an Mth-order tensor, the memory storage required will be $\sum_{m=1}^{M}(J_{m}\times r_{m})+\prod_{m=1}^{M}r_{m}$. Thus, it is clear that the dominant memory cost arises from the core tensor.
On the other hand, it can be shown that the sum of the squares of the elements of a tensor is equal to the sum of the squares of the elements of the core tensor, expressed as follows:
@@
$=\left\|\mathcal{Z}\right\|^{2}-2\left<\mathcal{Z}\times_{1}\mathbf{U}^{(1)t}\times_{2}\mathbf{U}^{(2)t}\dots\times_{m}\mathbf{U}^{(m)t}\dots\times_{M}\mathbf{U}^{(M)t},\mathcal{G}\right>+\left\|\mathcal{G}\right\|^{2}$
$=\left\|\mathcal{Z}\right\|^{2}-2\left<\mathcal{G},\mathcal{G}\right>^{2}+\left\|\mathcal{G}\right\|^{2}$
$=\left\|\mathcal{Z}\right\|^{2}-\left\|\mathcal{G}\right\|^{2}.$
It is also demonstrated that the sum of the squares of the core tensor is concentrated in a relatively small number of elements. Based on these two considerations, it may be reasonable to consider and save only the elements of the core that have the highest squared values. Moreover, instead of saving a fixed number of columns for each component matrix, i.e., $\mathbf{U}(J_{m}\times r_{m})$ for $m=1,\dots,M$, and a core tensor with corresponding dimensions, it is possible to save the full component matrices, i.e., $\mathbf{U}(J_{m}\times J_{m})$ for $m=1,\dots,M$, and only the elements of the core tensor with the highest magnitudes, ensuring that the ratio of the collected sum of squares to the total sum of squares is as high as desired.
@@
According to several strategies, such as compression ratio or the percentage of original information to retain, the quality measure of the original object is determined in Phase 1. In Phase 2, the order and dimensions of the new object must be established, where the order is higher than that of the original tensor, and the dimensions for each mode are lower. Heuristically, we have observed that tenSVD is more efficient when the new tensor approaches the shape of a hypercube.
In Phase 3, component matrices and the core tensor are calculated using SVD. Elements of the core tensor with the highest magnitudes are selected to ensure the desired level of quality in Phase 4. In Phase 5, the component matrices and selected elements, along with their relative positions in the core tensor, are saved in $\mathbf{U}^{(m)}$ ($m=1\dots M$) and $\mathbf{c}$, respectively. The data saved in this phase could be stored or transmitted to conserve space (in the case of backup) and bandwidth (in the case of transmission). In the final phase, the saved information is used to reconstruct the approximated original image.
@@
The steps of tenSVD are detailed in Algorithm 2, where it is possible to observe that both accuracy (given by Relative Error $\epsilon)$ and compression (given by compression ratio com) can be fixed.
Algorithm 2: tenSVD
Require: Tensor $\mathcal{X}(I_{1},I_{2},\dots,I_{N})$; $\epsilon$ (or com)
Reshape $\mathcal{X}\Rightarrow\mathcal{Z}(J_{1},J_{2},\dots,J_{J})$,
        with $J_{1}\cong J_{2}\cong\dots\cong J_{M}\cong M$ and $\prod_{i=1}^{N}I_{i}=\prod_{m=1}^{M}J_{m}$
for $m$ in $1,\dots,M$ do
        Consider the unfolding $\mathbf{Z}_{(m)}$ of $\mathcal{Z}$
        $\mathbf{U}^{(m)}\leftarrow$ $R_{m}$ left singular vectors of SVD($\mathbf{Z}_{(m)}\mathbf{Z}_{(m)}^{t}$)
end for
$dim\leftarrow size\left(\mathbf{U}^{(1)}\right)+\dots+size\left(\mathbf{U}^{(M)}\right)$
$\mathcal{G}\leftarrow\mathcal{Z}\times_{1}\mathbf{U}^{(1)t}\times_{2}\mathbf{U}^{(2)t}\times_{3}\dots\times_{M}\mathbf{U}^{(M)t}$
Select element of $\mathcal{G}$
$\mathbf{g}\leftarrow vec(abs(\mathcal{G}))$
Order in decreasing way the elements of $\mathbf{g}$
$r\leftarrow 1$
$\mathbf{\hat{g}}\leftarrow 0$
while $\left\|sum(\mathbf{g})-sum(\mathbf{\hat{g}})\right\|/\left\|sum(\mathbf{g})\right\|\geq\epsilon$ (or $dim/size\left(\mathcal{Z}\right)\leq com$) do
       $\mathbf{\hat{g}}\leftarrow\mathbf{\hat{g}}+\mathbf{g}[r]$
       $r\leftarrow 1+1$
        Set in $\mathbf{c}$ the elements of $\mathcal{G}$ selected and their relative position
        $dim\leftarrow dim+size(\mathbf{c})$
end while
return $\mathbf{c}$, $\mathbf{U}^{(1)}\dots\mathbf{U}^{(M)}$, $I_{1}\dots I_{N}$
end
In the first stage of the procedure, the elements are rearranged into an M-order tensor with dimensions $J_{1},J_{2},\dots,J_{M}$. The total number of elements, i.e., $\prod_{i=1}^{N}I_{i}$, is divided into prime factors. The smaller values are aggregated in increasing order until the minimum of them is approximately equal to M.
In the second stage, the HOSVD algorithm is applied to the new tensor $\mathcal{Z}$. The full left singular vectors and core tensor are calculated. From a computational point of view, it is more efficient to work on $\mathbf{Z}_{(m)}\mathbf{Z}_{(m)}^{t}$ than on the m-mode unfolding $\mathbf{Z}_{(m)}$.
In the final stage, the highest absolute values and their positions in the core tensor are selected as a vector, while ensuring that the fixed relative error or level of compression is achieved. Finally, these elements and component matrices are saved.
@@ 4 Case Study
The tenSVD algorithm was implemented in R to evaluate its performance for lossy compression. The algorithm was compared with the HOSVD algorithm (using the ‘tensr’ package). The analysis was conducted on an Intel(R) Xeon(R) Gold 6288 CPU running at 2.10 GHz (16 cores) with 512 GB of RAM, and using R version 4.2.0.
Three images — the Veiled Christ, the Doryphoros of Polykleitos, and the Royal Palace of Caserta — are shown in Fig.2. These images were converted into datasets: Fig.2 (a) has 880 rows and 1,240 columns, Fig.2 (b) has 1,792 rows and 2,560 columns, and Fig.2 (c) has 7,200 rows and 5,400 columns, all represented with three channels (RGB). Thus, three third-order tensors with dimensions $(880\times 1,240\times 3)$, $(1792\times 2560\times 3)$, and $(7200\times 5400\times 3)$ were generated.
Figure 2: Veilled Christ (a), Royal Palace of Caserta (b), Dorophos (c)
Additionally, a short video of the Athletics 100 meters final at the Paris 2024 Olympic Games was acquired. The grDevices package (version 4.4.0) was used to convert the three images into tensors. The Av package (version 0.9) was used to split the video into PNG images at a rate of three images per second. These frames were then converted into a fourth tensor with dimensions $(30\times 360\times 640\times 3)$.
@@ 4.1 Evaluation Metrics
Mean Squared Error (MSE), Relative Error (ERR ) and Peak-Signal-to-Noise Ratio (PSNR) are widely used to measure the accuracy of Tucker decomposition, while PSNR specifically assesses the quality of compressed images or videos.
@@
Moreover, to measure compression effectiveness in term of space and time, two additional metrics are used:
@@
    Compression Ratio (CR) is defined as the ratio between the size of the original tensor and the number of elements in the compression representation (storage cost).
    CR=compressed  data  sizeoriginal  data  $\textrm{CR}=\frac{\textrm{compressed \ data \ size}}{\textrm{original \ data \ size}}$
@@
    Processing time (TIME) refers to the actual elapsed time from the start of the algorithm execution.
@@ 4.2 Results
To compress the images, three different sets of parameters were chosen for HOSVD: $r_{1}=100$, $r_{2}=100$, $r_{3}=3$ (a); $r_{1}=200$, $r_{2}=200$, $r_{3}=3$ (b); and $r_{1}=600$, $r_{2}=600$, $r_{3}=3$ (c). The CR values observed from HOSVD were used as parameters for tenSVD to obtain comparable results in terms of memory cost.
In Figures 4, 5, and 6, the reconstructed images are displayed, with HOSVD results shown on the upper side and tenSVD results on the bottom side. The different scenarios demonstrate that both algorithms provide the same accuracy in terms of quality.
Figure 4: HOSVD CR$=93.7\%$ (a), CR$=83.4\%$ (b), and CR$=28.2\%$ (c); tenSVD CR$=93.7\%$ (d), CR$=83.4\%$ (e), and CR$=28.2\%$ (f).
Figure 5: HOSVD CR$=96.6\%$ (a), CR$=92.8\%$ (b), and CR$=83.2\%$ (c); tenSVD CR$=96.6\%$ (d), CR$=92.8\%$ (e), and CR$=83.2\%$ (f).
Figure 6: HOSVD CR$=98.9\%$ (a), CR$=97.7\%$ (b), and CR$=92.5\%$ (c); tenSVD CR$=98.9\%$ (d), CR$=97.7\%$ (e), and CR$=92.5\%$ (f).
In Table 1, all diagnostic measures are summarized. It is clear that all scenarios show very comparable values in terms of accuracy and quality, as demonstrated in Figures 4, 5, and 6. However, with fixed compression ratios, it is evident that tenSVD consistently provides better efficiency. Specifically, it is four times faster than HOSVD for the Veiled Christ and 26 times faster for the Doryphoros of Polykleitos, with both results based on scenario (c).
@@
Table 1: Performance comparison of HOSVD (H) and tenSVD (T) for the Veilled Christ, the Royal Palace of Caserta and the Doryphoros of Polykleitos with different compression ratio scenarios.
Regarding the Olympic Games Paris 2024 videos, a fixed set of parameters was analyzed: $r_{1}=30,\ r_{2}=300,\ r_{3}=300,\ r_{4}=3$. The HOSVD yielded the following metrics: MSE$=0.0013$, ERR$=0.0594$, PSNR$=28$, CR$=41\%$. In contrast, tenSVD analyzed the same compression ratio, producing comparable metrics: MSE$=0.0036$, ERR$=0.0952$, PSNR$=24$The tenSVD algorithm was approximately 2.5 times faster than HOSVD (22.024 seconds vs. 54.403 seconds). In Figure 7 presents the original frames 1, 10, 15, 20, and 30 (on the left side), the reconstructed frames after HOSVD (in the center), and the reconstructed frames after applying tenSVD (on the right side).
@@ 5 Conclusion and Findings
For the same level of quality and compression, it has been demonstrated that tenSVD is a faster algorithm compared to HOSVD. Moreover, the results indicate that the speed improvement can be quite significant in several scenarios. In Table 1, it can be observed that, on average, tenSVD is about 5.8 times faster than HOSVD for the 3,273,600 elements of the $880\times 1,240\times 3$ tensor, 10.1 times faster for the 13,762,560 elements of the $1,792\times 2,560\times 3$ tensor, and 27.9 times faster for the 32,531,040 elements of the $2,378\times 4,560\times 3$ tensor. In the case of the $30\times 360\times 640\times 3$ tensor, it is 2.5 times faster. Thus, the difference in computational time depends more on the shape of the tensor than on its size.
To further verify this, a simulation study was conducted on eight different image standards: $1280\times 720\times 3$ (HD), $1920\times 1080\times 3$ (Full HD), $2048\times 1080\times 3$ (2K), $2560\times 1440\times 3$ (4K), $3840\times 2160\times 3$ (4K UHD), $5120\times 2880\times 3$ (5K), $6144\times 3456\times 3$ (6K), and $7680\times 4320\times 3$ (8K). Datasets were generated 25 times using a uniform distribution. The times were measured in seconds, and summary statistics are given in Appendix A.
Figure 8 illustrates the elapsed time ratio between the two algorithms, showing that tenSVD is consistently faster than HOSVD, with the best results noted for the EK scenario.
Figure 8: Elapsed time ratio between HOSVD and tenSVD for scenario $1280\times 720\times 3$ (HD), $1920\times 1080\times 3$ (FullHD), $2048\times 1080\times 3$ (TwoK), $2560\times 1440\times 3$ (4HD), $3840\times 2160\times 3$ (QKUHD), $5120\times 2880\times 3$ (FK), $6144\times 3456\times 3$ (SK), and $7680\times 4320\times 3$ (EK)
@@
It appears that tenSVD is more efficient when it can reshape the original tensor into a form close to a hypercube. Other factors that influence efficiency include the initial shape of the tensor. In the case of higher quality images, the tensor is necessarily highly unbalanced because one mode has a dimension of 3 (RGB channels). In such cases, tenSVD demonstrates significant performance advantages in terms of computational cost.
Unfortunately, tenSVD is not suitable for several other applications. In situations where data is generated by latent variables, it cannot effectively analyze the underlying latent variables that caused the data generation, as the original structure is lost when the tensor is reshaped into a higher-order form. If the total core energy does not change when the tensor is reshaped, it is not possible to obtain the true orthonormal component matrices. In this context, HOSVD remains effective, but its accuracy may be compromised in the presence of significant noise.
Iterative algorithms such as Higher Order Orthogonal Iteration (HOOI) [9, 8], Sequentially Truncated Higher-Order Singular Value Decomposition (stHOSVD) [14], and Gradient Descent methods generally yield higher accuracy [1]. These algorithms have the advantage of converging to a better fit and are generally versatile, allowing for adaptation to many different scenarios and constraints. However, their iterative nature may lead to convergence issues, possibly getting stuck in local minima, and they could be very expensive in terms of both time and energy consumption.
ACKNOWLEDGMENTS
Data availability statement The data used in the paper are available either through the provided link with supplementary material or R packages.
Funding This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declarations
Conflict of interest The authors have no conflict of interest.
Ethical Approval Simulated data was used in the assessment of the proposed methods.
Open Access This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this licence, visit http://creativecommons.org/licenses/by/4.0/.

### v2

@@ 3.2 Memory cost
When lossy compression is the strategy, to reduce the memory cost the ($r_{1},\ r_{2},\dots,\ r_{N}$) are fixed lower than ($R_{1},\ R_{2},\dots,\ R_{N}$). And in case the elements are reordered into an higher order tensor, it is not clear if the memory cost will be higher. Storage requirement are given by the sum of memory cost for each component matrix $\sum_{n=1}^{N}I_{n}\times r_{n}$, and core tensor $\prod_{n=1}^{N}r_{n}$. If the elements are rearranged into an $M$th order tensor, memory storage required will be $\sum_{m=1}^{M}(J_{m}\times r_{m})+\prod_{m=1}^{M}r_{m}$. Thus it is clear that the dominant memory cost is due to the core tensor.
On the other side, it is possible to show that the energy (in terms of sum of the squared of elements) of a tensor is equal to the energy of core tensor, as
@@
$\left\|\mathcal{Z}\right\|^{2}-2\left<\mathcal{Z}\times_{1}\mathbf{U}^{(1)t}\times_{2}\mathbf{U}^{(2)t}\dots\times_{m}\mathbf{U}^{(m)t}\dots\times_{M}\mathbf{U}^{(M)t},\mathcal{G}\right>+\left\|\mathcal{G}\right\|^{2}$
$\left\|\mathcal{Z}\right\|^{2}-2\left<\mathcal{G},\mathcal{G}\right>+\left\|\mathcal{G}\right\|^{2}$
$\left\|\mathcal{Z}\right\|^{2}-\left\|\mathcal{G}\right\|^{2}.$
In Appendix A, it is shown too that core tensor energy is concentrated in a relatively small number of elements. Based on these two considerations, it could be reasonable consider and save only the elements of core that have the highest square values. Moreover, instead to save a fixed number of columns for each component matrices, i.e. $\mathbf{U}$ $(J_{m}\times r_{m})$ for $m=1,\dots,M$, and a core tensor with correspondent dimension, it is possible to save the full component matrices, i.e. $\mathbf{U}$ $(J_{m}\times J_{m})$ for $m=1,\dots,M$, and only the elements of core tensor with the highest magnitude, so that the ratio of energy collected is as high as researched.
@@
According to several strategies, such as compression ratio or the percentage of original information to retain, the quality measure of the original object is fixed (Phase 1). In Phase 2, the order and dimensions of the new object must be decided, where the order is higher than that of the original tensor and the dimensions for each mode are lower. Heuristically, we have observed that tenSVD is more efficient when the new tensor is close to a hypercube. In Phase 3, component matrices and the core tensor are calculated using SVD. Elements of the core tensor with highest magnitude are selected to ensure the desired level of quality (Phase 4). In Phase 5, component matrices and selected elements with their relative positions in the core tensor are saved in $\mathbf{U}$ and $\mathbf{c}$, respectively. Data saved in this phase could be stored or transmitted to save space (in case of backup) and bandwidth (in case of transmission). In the last Phase, the saved information is used to reconstruct the approximated original image.
The tenSVD steps are detailed given in Algorithm 2, where it is possible to observe that both accuracy, given by Relative Error ($\epsilon$), or compression, given by compression ratio (com), can be fixed.
In first stage of the procedure, to rearranged the elements into an $M$-order tensor with dimensions $J_{1},J_{2},\dots,J_{M}$, total number of elements, i.e. $\prod_{i=1}^{N}I_{i}$, is divides in prime factors, ordered increasing way smaller values are aggregated until the minimum of them is about equal to $M$.
In second stage, HOSVD algorithm is applied on new tensor $\mathcal{Z}$. Full left singular vectors and core tensor is given. From computational point of view, it is more efficient to work on $\mathbf{Z}_{(m)}\mathbf{Z}_{(m)}^{t}$ then on the $m$-mode unfolding $\mathbf{Z}_{(m)}$.
[…]
@@ 4 Case study
tenSVD algorithm was implemented in R and, to evaluate the performance for lossy compression problem, it is compared with HOSVD algorithm (package ‘tensr’ and ’rrcov3way’). The analysis was conducted using an Intel(R) Xeon(R) Gold 6288 CPU 2.10GHz 2.10 GHz (16 cores) with 512GB RAM, and R version 4.2.0.
Three pictures representing the Veiled Christ, the Dorophos, and the Royal Palace of Caserta are shown in Fig. 2. They are converted in datasets, Fig. 2 (a) has 880 rows and 1,240 columns, Fig. 2 (b) has 1,792 rows and 2,560 columns, and Fig. 2 (c) has 7,200 rows and 5,400 columns, all with three channels (RGB). Thus, three third-order tensors with dimensions $(880\times 1,240\times 3)$, $1792\times 2560\times 3$, and $7200\times 5400\times 3$ were generated. Moreover, a short video of the Athletics 100 meters final of the Olympic Games Paris 2024 has been acquired. Package grDevices 4.4.0 was used to convert the three pictures in tensors. Av package 0.9 was used to split the video in png images, where the rate was fixed in three images per second. These frames were converted in a fourth tensor with dimension $(30\times 360\times 640\times 3)$.
Figure 2: The Veilled Christ (a),the Royal Palace of Caserta (b), the Dorophos (c)
@@ 4.1 Evaluation Metrics
Mean Squared Error (MSE), Relative Error (ERR ) and Peak-Signal-to-Noise Ratio (PSNR) are widely used to measure the accuracy of Tucker decomposition, while PSNR is used to measure the quality of compressed images or videos. And they are defined as follows
@@
Figure 3: Used measurement protocol
Moreover, to measure compression effectiveness in term of space and time, two other metrics are used:
@@
    Compression Ratio (CR) is the ratio between the size of the original tensor and the number of elements of the compression representation (storage cost).
    $\textrm{CR}=\frac{\textrm{compressed \ data \ size}}{\textrm{original \ data \ size}}$
@@
    Processing time (TIME) is the real elapsed time since the algorithm was started.
@@ 4.2 Results
To compress the images, three different sets of parameters are chosen for HOSVD, $r_{1}=100$, $r_{2}=100$, $r_{3}=3$ (a), $r_{1}=200$, $r_{2}=200$, $r_{3}=3$ (b), $r_{1}=600$, $r_{2}=600$, $r_{3}=3$ (c), and CR values observed by t-HOSVD are used as parameters into tenSVD in order to get comparable results in term of memory cost.
In Figures 45, and 6 are shown the reconstructed images with HOSVD (upper side) and tenSVD (bottom side). The different scenarios show that both algorithms ensure the same accuracy in term of quality.
[…]
In Table 1 all diagnostic measures are summarized. It is clear that all scenarios show very comparable values in term of accuracy and quality as well as showed into Figures 4 5 6, but with fixed compression ratios, it is possible to observe that tenSVD assure always best results in term of efficiency. It is four 4 times faster than t-HOSVD for the Veilled Christ and 26 times faster for the Dorophos, both results is given for scenario (c).
@@
Table 1: Performance comparison of HOSVD (H) and tenSVD (T) for Veilled Christ, Royal Palace of Caserta and Dorophos with different compression ratio scenarios.
[…]
@@ 5 Conclusion and findings
For the same level of quality and compression, it is shown that tenSVD is a faster algorithm than HOSVD. Moreover, it was possible to observe that results could be impressively faster in several scenarios. In Table 1 it is possible to observe that in average tenSVD is about 5.8 times faster than HOSVD for the 3,273,600 elements of $880\times 1,240\times 3$ tensor, 10.1 times for the 13,762,560 elements of $1,792\times 2,560\times 3$ tensor, and 27.9 times for the 32,531,040 elements of $2,378\times 4,560\times 3$ tensor. When it is only 2.5 times faster for the 20,736,000 elements $30\times 360\times 640\times 3$ tensor. Thus, the difference of computational time depend from the shape more than size of tensor.
To verify this, a simulation study was run on 8 different image standards: $1280\times 720\times 3$(HD), $1920\times 1080\times 3$ (FullHD), $2048\times 1080\times 3$ (TwoK), $2560\times 1440\times 3$ (FourHD), $3840\times 2160\times 3$ (FourKUHD), $5120\times 2880\times 3$ (FiveK), $6144\times 3456\times 3$ (SixK), $7680\times 4320\times 3$ (EightK). Datasets were generated 25 times by uniform distribution. Time was measured in seconds and summary statistics are given in Table 1, when in Figure 8 shows the elapsed time ratio between the two algorithms. tenSVD is always faster than HOSVD and the best results is for EightK scenario.
Figure 8: Computational time ratio (HOSVD/tenSVD) across different image standards
@@
Our heuristic, it seems that tenSVD is higher efficient when it can reshape original tensor in close to hypercube. Other elements that influence the efficiency is starting shape of tensor. In case of higher quality images, tensor is necessary high unbalanced, because one mode has dimension 3 (RGB channels). In these cases tenSVD gives a huge performance in term of computational cost.
Unfortunately, tenSVD is not suitable in several other applications. It cannot find the true latent variables that generate data, because original structure is missed when tensor is reshaped in an higher-order ones. If the total of core energy doesn’t change when tensor is reshaped, it is not possible to get the true orthonormal component matrices. In this contest, HOSVD works but it accuracy could be low in presence of significant noise. Iterative algorithms as Higher-Order Orthogonal Iteration (HOOI) [8, 3], Sequentially Truncated Higher-Order Singular Value Decomposition (stHOSVD) [4] or Gradient Descent methods generally yields higher accuracy [14] [12]. These algorithms have advantage to converge at better fit and generally they are versatile and can be adapted to many different scenarios and constraints. On the other side, their iterative nature might get stuck in local minima and could be very expensive in term of time consuming and relative sustainability in term of energy required.